let F^mxn be the set of all mxn matrices over the field F. is F^mxn is a vector space?


1
Expert's answer
2021-06-29T03:41:52-0400

Let V be the vector space 


Let "V={A=(a_{ij})\in M_{m\times n}: a_{ij}\in R}"


Let "X=(x_{ij}),Y=(y_{ij})\in V" and "\alpha,\beta\in Q"


Now lets us check if vector "\alpha X+\beta Y" is an element of V.


"\alpha X+\beta Y=\alpha(x_{ij})+\beta (y_{ij})"


        "=(\alpha x_{ij})+(\beta y_{ij})\\[9pt]

 =(\alpha x_{ij}+\beta y_{ij})"


Since "(\alpha x_{ij})+(\beta y_{ij})\in R \forall i" , Vector "\alpha X+\beta Y" is an element of V.


So V is the vector space over the field F. Which contains all the "m\times n" matrices.